# Weakly Supervised Learning for Facial Behavior Analysis: A Review

### v3

@@ 4.2.3 Discussion
4.3 Experimental Evaluations
4.3.1 Datasets
To cover the wide range of facial expressions, several databases have been developed with intensity levels of expressions such as pain and AUs. In this section, we describe a dataset for regression that does not appear in Section 3.4.1.
@@ 4.3.2 Experimental Protocol
4.4 Critical Analysis and Results
@@
The intensity estimation of facial expressions or AUs is more challenging than the task of classification due to the complexity of capturing the subtle variations of facial appearance. Therefore, the task of intensity estimation remains an under-researched problem compared to the task of classification. To the best of our knowledge, there are no works on inaccurate annotations with intensity estimation despite the high level of uncertainty in intensity labels. Since temporal dynamics plays a crucial role in conveying significant information for the task of estimating the intensity level, [89] and [88] modeled the relevant ordinal relationship across temporal frames by incorporating intensity and relevance smoothness into the objective function. However, both of these approaches rely on classical ML models, which have been outperformed by Praveen et al. [84, 14] that leverage 3D-CNN models as shown in Table IV. The task of estimating intensity is relatively less explored in the context of inexact annotations when compared to incomplete annotations.
@@
For incomplete annotations, [91, 94] and [92] explored temporal, feature, and label smoothness in addition to the temporal relevance, where the former shows better performance by exploring more constraints. [93] explored attention mechanisms at both the feature and label levels, showing an improvement over prior approaches [91, 94, 92]. [96] and [95] further improved the performance by leveraging large-scale pre-trained models based on self-supervised learning and RBMs to model AU relationships, where the latter has shown better performance. Compared to classical models [89, 88], deep models [95, 96] have shown significant improvements for both inexact and incomplete annotations. However, a challenge in using DL models for estimating the intensity of facial expressions or AUs is the requirement of a large number of representative samples with intensity annotations, which is costly and requires domain expertise. Estimation of intensity levels using DL models is therefore explored in the context of incomplete annotations by leveraging large-scale pre-trained models [96] or large-scale unlabeled data [94]. Similar to that of classification, DL models with inexact annotations still remain an under-researched problem for regression.
5 Open Problems and Opportunities
@@
This section describes the challenges for FER in the context of WSL. Although developing a robust FER system raises many challenges such as identity bias and data sparsity [5, 100], we emphasize the challenges and research directions relevant in the context of weak annotations.
5.1 Open Problems
@@
Training with Limited Reliable Annotations. As the data capture conditions change from controlled laboratory environments to uncontrolled real-world environments, the visual appearance of facial expressions exhibits a wide range of variations between people, depending on age, civilization, ethnicity, cosmetics, eyeglasses, etc. Moreover, facial expressions are sparse in videos, as they are often expressed in a subset of the frames along with neural frames. The problem is accentuated by ambiguous annotations, as they further reduce the number of training samples with accurate labels. In most machine learning-based applications, the performance of the system is highly influenced by the quantity of representative data, which has been reflected in the high level of accuracy obtained with DL models. Given the limited number of training samples with reliable annotations, the potential of DL models is not fully explored due to the overfitting and generalization capability. In the case of incomplete and inaccurate annotations, the problem of limited samples is circumvented by an abundance of unlabeled data [45, 48, 43] and noisy annotations obtained from the wild [61, 59, 67] respectively.
@@
Localizing AUs with Weak Image-Level Annotations. Localizing AUs with image-level AU annotations can be formulated in the context of inexact annotations. Some approaches have recently been proposed for localizing AU regions based on facial landmarks [101, 102, 103]. However, these approaches rely on prior knowledge of predefined AU attentions, which restrict the capacity of predefined AUs, and fail to capture the wide range of non-rigid AUs. Given the potential of local features related to AU and the tedious process of obtaining localized AU regions, there is a need to formulate the problem of localizing AU patches without predefined AU attention. To our knowledge, only two approaches have addressed the problem of AU localization without any prior information on pre-defined AU attention. [104] integrated the relations among AUs with an attention mechanism in an end-to-end DL framework by extracting multiscale features, while [105] explored self-supervised learning for localizing AUs using optical flow and further deployed in a cross-domain setting.
@@
Data Bias with Incomplete Annotations. Leveraging large-scale face recognition (FR) datasets is a promising direction to address the limitations of labeled FER datasets. However, leveraging FR datasets as unlabeled data poses the problem of a distribution mismatch between small-scale FER datasets and large-scale unlabeled FR datasets, which can deteriorate system performance. To the best of our knowledge, only two works attempted to solve this issue. [106] addressed the problem of class imbalance along with the mismatch of data distributions using base and adaptation networks in an iterative feedback mechanism, while [48] explored self-supervised learning with contrastive learning. Therefore, dealing with data set bias to cope with FER data with limited annotation remains a challenging open problem, which can be further explored to improve system performance.
@@
AU Recognition with Noisy Annotations. Most approaches for AU recognition have been explored in the context of inexact annotation (using expression labels) [36, 37, 38] or incomplete annotations (missing or incomplete labels) [50, 51, 54]. Manual annotation of AUs is more challenging compared to categorical expressions because of the increased range of facial behavior. Moreover, the annotation of AUs requires domain expertise certified by the FACS coding system, which is laborious, and thereby prone to annotation errors. Given the complex annotation process for AUs, most of the existing AU datasets are constrained by the number of coded AU samples. Although the number of facial images on the web has been growing at an exponential rate, leveraging large-scale datasets captured in the wild for AU recognition is relatively unexplored in the literature. Most current approaches for AU recognition with inexact or incomplete annotations rely on classical facial landmarks based on geometric features [36, 38]. Though few works explored RBMs for AU recognition [37], leveraging DL models for AU recognition with weak annotations requires a large-scale dataset. Therefore, there is a scope for research on AU recognition with inaccurate annotations using large in-the-wild datasets. Only two works [68, 69] (see Sections 3.3.2) have been explored in the literature for AU recognition with inaccurate annotations, which leaves much room for further improvement.
@@
Dimensional Affect Model with Weak Annotations. As described in Section 3.3, the problem of FER with inaccurate annotations is explored mainly in the context of classification. Although the problem of noisy annotations is more pronounced in the case of dimensional or ordinal annotations due to subtle variations between consecutive frames, few works have addressed the problem of inaccurate annotations in the dimensional model [107, 108]. Given the complexity of obtaining annotations in the dimensional model and the lack of techniques to handle noisy dimensional annotations, most of the datasets are developed for the task of classification of facial expressions or AUs, which is mentioned in Section 3.4.1. Though few datasets have been explored for the task of ordinal regression (as described in Section 4.3.1), the development of datasets for continuous dimensional models is rarely explored. To our knowledge, only two datasets, i.e. [11] and [9] have been explored for FER in a continuous dimensional model, however, several multimodal datasets are available [109, 110, 111]. Ordinal annotations of DISFA [75] are obtained by two FACS-certified experts and noisy annotations are reduced by evaluating the correlation between their annotations. For UNBC-McMaster [72], the ordinal annotations are obtained from three FACS coders, which were then reviewed by a fourth FACS coder and validated using Ekman-Friesen formulae [112]. Similarly, [9] relies on 12 expert annotators, and the annotations were further reviewed by two independent annotators. [11] obtained annotations from six trained experts and were doubly reviewed by two more annotators. Finally, the final labels are considered as the mean of the annotations for each sample. They have also conducted statistical analysis to evaluate the inter-annotator correlations. Therefore, modeling the dimensional labels remains a highly challenging open problem due to the high level of ambiguity in the annotations.
5.2 Potential Research Directions
This section presents potential research directions for the advancement of WSL methods in facial behavior analysis.
@@ 5.2 Potential Research Directions
Domain Adaptation. Domain Adaptation (DA) for FER is a promising line of research to deal with the problem of limited samples with reliable annotations. Adapting a model using unlabeled videos from relevant external datasets to counterfeit the problem of overfitting deep models. It also allows adapting the model to the target operational environment. It has been widely used for many applications related to facial analysis such as face recognition [113], facial expression recognition [84], etc. Similarly, domain generalization (DG) methods can also be exploited to improve the robustness of source models (without target data), especially for WSL with inexact or incomplete WSL scenarios. One example in this direction is the work of Wang et al. [114], who proposed an unsupervised DA approach for small target datasets using GANs, where GAN-generated samples are used to fine-tune the model pre-trained on the source dataset. Another work done by Zhu et al. [115] explored an unsupervised DA approach in the feature space, where the mismatch between the feature distributions of the source and target domains is minimized while improving the discrimination among the face images related to facial expressions. Shao et al. [116] exploited the collection of constrained images from the source domain with both AU and landmark labels, and an unpaired collection of unconstrained wild images from the target domain with only landmark labels. The features extracted from source and target images are disentangled to shape and text features, and the shape features (AU label information) of source images are fused with the texture information of the target feature. Therefore, the performance of FER can be enhanced using DA along with DL to handle data with limited annotations still harnessing the potential of deep networks.
@@
Deep Learning Models. Given their success, many researchers have leveraged DL models for various computer vision applications such as object detection, face recognition, etc., and showed significant improvement in performance over the traditional approaches in real-world conditions. However, the performance of deep models is not fully explored in the context of facial behavior analysis due to the limited training data and laborious task of annotations which demands human expertise. Despite the complex process of obtaining annotations for facial behavior, most of the existing approaches to FER based on DL have been focused on fully supervised settings. Shan et al. [12] provided a comprehensive survey on DL approaches for FER in the framework of supervised learning, and provided insight into the advantages and limitations of deploying deep models for FER. Recently, DL models have been explored in the context of incomplete and inaccurate annotations by leveraging the unlabeled and noisy wild datasets [43, 44, 59] for expression recognition. However, it is still at a rudimentary level for the recognition of AUs. Most of the approaches for AU recognition used geometric features based on facial landmark [41, 38, 37] or classical features such as LBP [40], SIFT [36], etc. In the case of regression tasks, very few works explored deep models by leveraging large-scale pre-trained models [96] or CNN with limited hidden layers [91]. [37] and [41] have explored RBMs to model the domain knowledge of dependencies among AUs and the relationship between expressions and AUs. For AU intensity estimation, two works [91, 94] used CNN with 3 layers, and two more works [96, 93] used Resnet-18, where [96] explored large-scale pre-trained models with self-supervised learning. To the best of our knowledge, no work has been done on expression detection with inexact annotations using deep models.
@@
Temporal Dynamics. In most approaches for FER in WSL, only short-term dynamics across the temporal frames are exploited. [117] and [118] used LBP-TOP features [18] while [31], [28], and [119] used frame aggregation – the maximum of feature vectors of the frames for modeling temporal dynamics. [33] capture the temporal order of the templates of the sequence by appending frame-level features of the sequence, while [91] captures temporal dynamics using contrastive appearance difference, i.e., the difference between apex frame and neutral frame. Recently, several DL models have been proposed to leverage the temporal dynamics of facial expressions over video sequences [120, 121]. Convolutional 3D (C3D) models [122] are also gaining much attention in computer vision for modeling spatio-temporal information across frames. Compared to a 2D-CNN with RNN, C3D is efficient in capturing short-term temporal information. Though LSTM and C3D techniques are widely explored for FER in fully supervised settings [123], [124], it is not yet effectively explored for the framework of WSL for FER. Therefore, LSTM and C3D techniques when deployed in a WSL setting for FER are expected to further enhance the performance of existing state-of-the-art approaches. Leveraging transformers to model the temporal dynamics by exploiting self-supervised learning to mitigate the problem of limited annotations is also expected to improve the performance of FER in WSL [125].
@@
Modeling Perception Uncertainty. A challenge that is relevant for FER with weak annotations is modeling perception uncertainty to closely reflect human perception of emotions. Most approaches in the literature rely on the assumption that annotations obtained from crowd-sourcing will be more accurate, yet this assumption is not verified in realistic scenarios. Even when multiple annotators are engaged, producing multiple diverse annotations, label uncertainty still remains in the emotion target as the number of annotators cannot be infinite. This inevitably introduces noise in annotations, which may degrade system performance. Motivated by this observation, [126] introduced the idea of modeling the uncertainty of emotion perception using soft labels instead of hard labels. Although the problem of modeling the perception uncertainty is unexplored in the literature, there is great potential for further exploration.
@@
Continuous Affect Models. The continuous dimensional model conveys a wider range of expressions than ordinal regression and plays a crucial role in capturing the subtle changes and context sensitivity of emotions. Compared to the task of classification and ordinal regression, WSL-based approaches for facial expressions or AUs are hardly explored in the context of continuous dimensional space though few endeavors have been made in the context of fully supervised learning [127], [128]. Hatice et al. [7] investigated the potential of a continuous dimensional model, and provided insight on state-of-the-art approaches and challenges associated with automatic continuous analysis and synthesis of emotional behavior. Given the intricate and error-prone process of obtaining annotations, there is an imperative need to formulate the problem of the continuous dimensional model in the framework of WSL to handle noisy annotations and alleviate the negative impact of unreliable annotations. Huang et al. [129] investigated the impact of annotation delay compensation and other post-processing operations for continuous emotion prediction of multi-modal data. As far as we know, there is only one work [130] based on WSL for the prediction of a continuous dimensional model, using valence and arousal in a multi-modal framework. They have reduced the noise of unreliable labels by introducing temporal label, which incorporates contextual information by considering the labels within a temporal window for every time step. They have further used a robust loss function that ignores small errors between predictions and labels in order to further reduce the impact of noisy labels. Therefore, there is much room for improvement to augment the performance of the FER system in the continuous dimensional model using WSL-based approaches.
@@
Multimodal Learning. Multimodal analysis has gained much attention in recent years to enhance the overall accuracy and robustness of FR systems over uni-modal approaches. Exploring multiple modalities may provide more diverse and comprehensive information, which can help reduce the cost of annotation, as well as annotation error or low-quality annotated datasets [131]. Inspired by their performance, several data sets are developed to validate multimodal approaches under challenging real-world scenarios [109], [132]. To encourage progress in multimodal emotion recognition, several datasets including audio, visual, text, etc. have been introduced [109, 11, 110, 133]. Several works have been done based on audio-visual fusion for emotion recognition [134, 135, 136, 137] in the framework of supervised learning. To our knowledge, only one work [130] has been published on WSL approaches for multimodal affect recognition using audio and visual features. Inspired by the invariance of thermal images to illumination, some approaches have been proposed to exploit thermal images with RGB images in a complementary fashion to augment the performance of the FER system [138, 139]. Another line of research is dealing with noisy data caused by facial pose or occlusion, which may lead to ambiguous annotations. To overcome the problem of pose-variance and occlusion, 3D data has been explored to obtain the comprehensive information displayed by the face and capture the subtle changes of facial AUs in detail using the depth of the facial surface. For instance, AU18 (lip pucker) is hard to differentiate from AU10+AU17+AU24 in a 2D frontal view. Sandach et al. [140] provided a comprehensive survey on datasets and FER systems relevant to 3D or 4D data. Inspired by the performance of optical flow in action recognition, a few approaches have been proposed to capture facial muscular movements using optical flow as it converts the sequence of motion information to static textures [141, 142]. Benjamin et al. [143] investigated the effectiveness of optical flow in recognizing full expressions, as well as micro-expressions, in near-frontal images. As far we know, only one work [144] has explored optical flow for FER with minimal annotations i.e., sequence-level labels.
@@
Auxiliary Spaces for label Uncertainty Exploring the auxiliary label space of relevant tasks has recently gained much attention for dealing with label uncertainty. The idea is to leverage the information available in the auxiliary label space of related but distinct tasks to refine the label uncertainty of the main FER task. Chen et al. [145] explored the labels of AU recognition and facial landmark detection to effectively learn the label distributions based on the assumption that neighboring facial images in the auxiliary label space should have similar expression distributions. Le et al. [146] showed that leveraging the neighborhood information in the valence-arousal space helps to obtain more semantic label distributions, and further improves system performance. Unlike the above two approaches, Liu et al. [147] exploited the auxiliary tasks of AU detection and valence-arousal measurement to learn to relabel uncertain samples and mitigate the class imbalance respectively. Therefore, leveraging the auxiliary label space is a promising line of research to deal with the problem of label uncertainty, and can also be extended to deal with limited annotations.
@@
Few Shot Learning Few-Shot Learning (FSL) is another emerging machine learning paradigm that deals with the problem of learning from limited annotated samples of novel classes. The basic idea of FSL is to generalize a pre-trained model to new classes by learning from a limited set of labeled samples of these classes (support set) [148]. Although FSL does not exactly correspond to any of the WSL scenarios, it overlaps somewhat with the framework of incomplete annotations, as both cases deal with the problem of limited accurate annotations. Recently, FSL has also gained attention in FER as a solution to the problem of novel classes for compound emotion recognition [149, 125, 150]. Another line of research with FSL for FER is to adapt to new subjects as people express their emotions differently from each other [151]. Most of the existing approaches pertinent to FSL for FER have been evaluated in cross-domain settings, where the target domain may have few labeled samples of novel classes [149, 150, 125]. Xinyi et al. [149] proposed a novel emotion-guided similarity network to deal with the limited annotations of compound expressions, where the pre-trained model is generalized to limited samples of novel classes (support set) of compound emotions. Zou et al. [150] also addressed the problem of compound FER in a cross-domain setting using a cascaded decomposition network for FSL. Chen et al. [125] explored self-supervised vision transformers by jointly pertaining with multiple pretext tasks, and leveraged FSL to train the deep model with fewer labeled samples. Besides these few papers on compound emotion recognition, FSL has not been explored for AU recognition. Therefore, the use of FSL techniques is a promising line of research to deal with limited labeled samples for FER.
6 Conclusion

### v4

@@ IV-B3 Discussion
IV-C Experimental Evaluations
IV-C1 Datasets
To cover the wide range of facial expressions, several databases have been developed with intensity levels of expressions such as pain and AUs. In this section, we describe a dataset for regression that does not appear in Section III-D1.
@@ IV-C2 Experimental Protocol
IV-D Critical Analysis and Results
@@
The intensity estimation of facial expressions or AUs is more challenging than the task of classification due to the complexity of capturing subtle variations in facial appearance. Therefore, it remains an under-researched problem compared to the task of classification. To the best of our knowledge, there are no works on inaccurate annotations with intensity estimation despite the high level of uncertainty in intensity labels. Since temporal dynamics plays a crucial role in conveying significant information for the task of estimating the intensity level, [89] and [88] modeled the relevant ordinal relationships across temporal frames by incorporating intensity and relevance smoothness into the objective function. However, both of these approaches rely on classical ML models. They have been outperformed by [85, 14] that leverage 3D-CNN models as shown in Table IV. The task of estimating intensity is less explored in the context of inexact annotations when compared to incomplete annotations.
@@
For incomplete annotations, [92, 95] and [93] explored temporal, feature, and label smoothness in addition to the temporal relevance, where the former shows better performance by leveraging more constraints. [94] explored attention mechanisms at both the feature and label levels, showing an improvement over prior approaches [92, 95, 93]. [97] and [96] further improved performance by leveraging large-scale pre-trained models based on self-supervised learning and RBMs to model AU relationships, where the latter has shown better performance. Compared to classical ML models [89, 88], DL models [96, 97] have shown significant improvements for both inexact and incomplete annotations. However, a challenge in using DL models for estimating the intensity of facial expressions or AUs is the requirement of a large number of representative samples with intensity annotations, which is costly and requires domain expertise. Estimation of intensity levels using DL models is therefore explored in the context of incomplete annotations by leveraging large-scale pre-trained models [97] or large-scale unlabeled data [95]. Similar to that of classification, DL models with inexact annotations remain an under-researched problem for regression.
V Open Problems and Opportunities
@@
This section describes the challenges for FER in the context of WSL. Although developing a robust FER system raises many challenges such as identity bias and data sparsity [5, 101], we emphasize the challenges and research directions for FER in the context of weak annotations.
V-A Open Problems
@@
(a) Training with Limited Reliable Annotations. As the data capture conditions change from controlled laboratory environments to uncontrolled real-world environments, the visual appearance of facial expressions exhibits a wide range of variations between people, depending on age, civilization, ethnicity, cosmetics, eyeglasses, capture conditions, etc. Moreover, facial expressions may be sparse in videos, as they are often expressed in a subset of the frames along with neural frames. The problem is accentuated by ambiguous or uncertain annotations, as they further reduce the number of training samples with reliable labels. The performance of the ML/DL models is typically influenced by the quantity of representative data. Given a limited number of training samples with reliable annotations, the potential of DL models is not fully explored due to the overfitting and generalization capability. Though large-scale unlabeled or noisy annotated data are leveraged for the context of incomplete [47, 50, 45] and inaccurate annotations [63, 61, 69], exploring the potential of DL models with weak annotations for FER is not fully explored.
@@
(b) Localizing AUs with Weak Image-Level Annotations. Localizing AUs with image-level AU annotations can be formulated in the context of inexact annotations. Some approaches have recently been proposed for localizing AU regions based on facial landmarks [102, 103, 104]. However, these approaches rely on prior knowledge of predefined AU attentions, which restrict the capacity of predefined AUs, and fail to capture the wide range of non-rigid AUs. Given the potential of local features related to AU and the tedious process of obtaining localized AU regions, there is a need to formulate the problem of localizing AU patches without predefined AU attention. To our knowledge, only two approaches have addressed the problem of AU localization without any prior information on pre-defined AU attention. [105] integrated the relations among AUs with an attention mechanism in an end-to-end DL framework by extracting multiscale features, while [106] explored self-supervised learning for localizing AUs using optical flow and further deployed in a cross-domain setting.
@@
(c) Data Bias with Incomplete Annotations. Leveraging large-scale face recognition (FR) datasets is a promising direction to address the limitations of labeled FER datasets. However, leveraging FR datasets as unlabeled data poses the problem of a distribution mismatch between small-scale FER datasets and large-scale unlabeled FR datasets, which can deteriorate system performance. To the best of our knowledge, only two works attempted to solve this issue. [107] addressed the problem of class imbalance along with the mismatch of data distributions using base and adaptation networks in an iterative feedback mechanism, while [50] explored self-supervised learning with contrastive learning. Therefore, dealing with data set bias to cope with FER data with limited annotation remains a challenging open problem, which can be further explored to improve system performance.
@@
(d) AU Recognition with Noisy Annotations. Most approaches for AU recognition have been explored in the context of inexact annotation (using expression labels) [38, 39, 40] or incomplete annotations (missing or incomplete labels) [52, 53, 56]. Annotating AUs is more challenging compared to categorical expressions because of the increased range of facial behavior. Moreover, the annotation of AUs requires domain expertise certified by the FACS coding system, which is laborious, and thereby prone to annotation errors. Given the complex annotation process for AUs, most of the existing AU datasets are constrained by the number of coded AU samples. Although the number of facial images on the web has been growing at an exponential rate, leveraging large-scale datasets captured in the wild for AU recognition is relatively unexplored in the literature. Most current approaches for AU recognition with inexact or incomplete annotations rely on classical facial landmarks based on geometric features [38, 40]. Though few works explored RBMs for AU recognition [39], leveraging DL models for AU recognition with weak annotations requires a large-scale dataset. Therefore, there is a scope for research on AU recognition with inaccurate annotations using large in-the-wild datasets. Only two works [70, 71] (see Sections 3.3.2) have been explored in the literature for AU recognition with inaccurate annotations, which leaves much room for further improvement.
@@
(e) Dimensional Affect Model with Weak Annotations. As described in Section 3.3, the problem of FER with inaccurate annotations is explored mainly in the context of classification. Although the problem of noisy annotations is more pronounced in the case of dimensional or ordinal annotations due to subtle variations between consecutive frames, few works have addressed the problem of inaccurate annotations in the dimensional model [108, 109]. Given the complexity of obtaining annotations in the dimensional model and the lack of techniques to handle noisy dimensional annotations, most of the datasets are developed for the task of classification of facial expressions or AUs, which is mentioned in Section 3.4.1. Though few datasets have been explored for the task of ordinal regression (as described in Section 4.3.1), the development of datasets for continuous dimensional models is rarely explored. To our knowledge, only two datasets, i.e. [11] and [9] have been explored for FER in a continuous dimensional model, however, several multimodal datasets are available [110, 111, 112]. Ordinal annotations of DISFA [77] are obtained by two FACS-certified experts and noisy annotations are reduced by evaluating the correlation between their annotations. For UNBC-McMaster [74], the ordinal annotations are obtained from three FACS coders, which were then reviewed by a fourth FACS coder and validated using Ekman-Friesen formulae [113]. Similarly, [9] relies on 12 expert annotators, and the annotations were further reviewed by two independent annotators. [11] obtained annotations from six trained experts and were doubly reviewed by two more annotators. Finally, the final labels are considered as the mean of the annotations for each sample. They have also conducted statistical analysis to evaluate the inter-annotator correlations. Therefore, modeling the dimensional labels remains a highly challenging open problem due to the high level of ambiguity in the annotations.
V-B Potential Research Directions
@@
(a) Domain Adaptation. Unsupervised Domain Adaptation (UDA) for FER is a promising line of research to deal with the problem of limited samples with reliable annotations, by adapting a model using unlabeled videos from relevant external datasets to counterfeit the problem of overfitting deep models. It also allows adapting the model to the target operational environment. It has been widely used for many applications related to facial analysis such as face recognition [114], facial expression recognition [85, 115], etc. Similarly, domain generalization (DG) methods can also be exploited to improve the robustness of source models (without target data), especially for WSL with inexact or incomplete WSL scenarios. One example in this direction is the work of Wang et al. [116], who proposed an UDA approach for small target datasets using GANs, where GAN-generated samples are used to fine-tune the model pre-trained on the source dataset. Another work done by Zhu et al. [117] explored an UDA approach in the feature space, where the mismatch between the feature distributions of the source and target domains is minimized while improving the discrimination among the face images related to facial expressions. Shao et al. [118] exploited the collection of constrained images from the source domain with both AU and landmark labels, and an unpaired collection of unconstrained wild images from the target domain with only landmark labels. The features extracted from source and target images are disentangled to shape and text features, and the shape features (AU label information) of source images are fused with the texture information of the target feature. Therefore, the performance of FER can be enhanced using UDA along with DL to handle data with limited annotations still harnessing the potential of deep networks.
@@
(b) Deep Learning Models. Given their success, many researchers have leveraged DL models for various computer vision applications such as object detection, face recognition, etc., and showed significant improvement in performance over the traditional approaches in real-world conditions. However, the performance of deep models is not fully explored in the context of facial behavior analysis due to the limited training data and laborious task of annotations which demands human expertise. Despite the complex process of obtaining annotations for facial behavior, most of the existing approaches to FER based on DL have been focused on fully supervised settings. Li et al. [12] provided a comprehensive survey on DL approaches for FER in the framework of supervised learning, and provided insight into the advantages and limitations of deploying deep models for FER. Recently, DL models have been explored in the context of incomplete and inaccurate annotations by leveraging the unlabeled and noisy wild datasets [45, 46, 61] for expression recognition. However, it is still at a rudimentary level for the recognition of AUs. Most of the approaches for AU recognition used geometric features based on facial landmark [43, 40, 39] or classical features such as LBP [42], SIFT [38], etc. In the case of regression tasks, very few works explored deep models by leveraging large-scale pre-trained models [97] or CNN with limited hidden layers [92]. [39] and [43] have explored RBMs to model the domain knowledge of dependencies among AUs and the relationships between expressions and AUs. For AU intensity estimation, two works [92, 95] used CNN with 3 layers, and two more works [97, 94] used Resnet-18, where [97] explored large-scale pre-trained models with self-supervised learning. To the best of our knowledge, no work has been done on expression detection with inexact annotations using deep models.
@@
(c) Temporal Dynamics. In most approaches for FER in WSL, only short-term dynamics across the temporal frames are exploited. [119] and [120] used LBP-TOP features [18] while [33], [30], and [121] used frame aggregation – the maximum of feature vectors of the frames for modeling temporal dynamics. [35] capture the temporal order of the templates of the sequence by appending frame-level features of the sequence, while [92] captures temporal dynamics using contrastive appearance difference, i.e., the difference between apex frame and neutral frame. Recently, several DL models have been proposed to leverage the temporal dynamics of facial expressions over video sequences [122, 123]. Convolutional 3D (C3D) models [124] are also gaining much attention in computer vision for modeling spatio-temporal information across frames. Compared to a 2D-CNN with RNN, C3D is efficient in capturing short-term temporal information. Though LSTM and C3D techniques are widely explored for FER in fully supervised settings [125], [126], it is not yet effectively explored for the framework of WSL for FER. Therefore, LSTM and C3D techniques when deployed in a WSL setting for FER are expected to further enhance the performance of existing state-of-the-art approaches. Leveraging transformers to model the temporal dynamics by exploiting self-supervised learning to mitigate the problem of limited annotations is also expected to improve the performance of FER in WSL [127].
@@
(d) Modeling Perception Uncertainty. A challenge that is relevant for FER with weak annotations is modeling perception uncertainty to closely reflect human perception of emotions. Most approaches in the literature rely on the assumption that annotations obtained from crowd-sourcing will be more accurate, yet this assumption is not verified in realistic scenarios. Even when multiple annotators are engaged, producing multiple diverse annotations, label uncertainty still remains in the emotion target as the number of annotators cannot be infinite. This inevitably introduces noise in annotations, which may degrade system performance. Motivated by this observation, [128] introduced the idea of modeling the uncertainty of emotion perception using soft labels instead of hard labels. Although the problem of modeling the perception uncertainty is unexplored in the literature, there is great potential for further exploration.
@@
(e) Continuous Affect Models. The continuous dimensional model conveys a wider range of expressions than ordinal regression and plays a crucial role in capturing the subtle changes and context sensitivity of emotions. Compared to the task of classification and ordinal regression, WSL-based approaches for facial expressions or AUs are hardly explored in the context of continuous dimensional space though few endeavors have been made in the context of fully supervised learning [129], [130]. Hatice et al. [7] investigated the potential of a continuous dimensional model, and provided insight on state-of-the-art approaches and challenges associated with automatic continuous analysis and synthesis of emotional behavior. Given the intricate and error-prone process of obtaining annotations, there is an imperative need to formulate the problem of the continuous dimensional model in the framework of WSL to handle noisy annotations and alleviate the negative impact of unreliable annotations. Huang et al. [131] investigated the impact of annotation delay compensation and other post-processing operations for continuous emotion prediction of multi-modal data. As far as we know, there is only one work [132] based on WSL for the prediction of a continuous dimensional model, using valence and arousal in a multi-modal framework. They have reduced the noise of unreliable labels by introducing temporal label, which incorporates contextual information by considering the labels within a temporal window for every time step. They have further used a robust loss function that ignores small errors between predictions and labels in order to further reduce the impact of noisy labels. Therefore, there is much room for improvement to augment the performance of the FER system in the continuous dimensional model using WSL-based approaches.
@@
(f) Multimodal Learning. Multimodal analysis has gained much attention in recent years to enhance the overall accuracy and robustness of FR systems over uni-modal approaches. Exploring multiple modalities may provide more diverse and comprehensive information, which can help reduce the cost of annotation, as well as annotation error or low-quality annotated datasets [133]. Inspired by their performance, several data sets are developed to validate multimodal approaches under challenging real-world scenarios [110], [134]. To encourage progress in multimodal emotion recognition, several datasets including audio, visual, text, etc. have been introduced [110, 11, 111, 135]. Several works have been done based on audio-visual fusion for emotion recognition [136, 137, 138, 139] in the framework of supervised learning. To our knowledge, only one work [132] has been published on WSL approaches for multimodal affect recognition using audio and visual features. Inspired by the invariance of thermal images to illumination, some approaches have been proposed to exploit thermal images with RGB images in a complementary fashion to augment the performance of the FER system [140, 141]. Another line of research is dealing with noisy data caused by facial pose or occlusion, which may lead to ambiguous annotations. To overcome the problem of pose-variance and occlusion, 3D data has been explored to obtain the comprehensive information displayed by the face and capture the subtle changes of facial AUs in detail using the depth of the facial surface. For instance, AU18 (lip pucker) is hard to differentiate from AU10+AU17+AU24 in a 2D frontal view. Sandach et al. [142] provided a comprehensive survey on datasets and FER systems relevant to 3D or 4D data. Inspired by the performance of optical flow in action recognition, a few approaches have been proposed to capture facial muscular movements using optical flow as it converts the sequence of motion information to static textures [143, 144]. Benjamin et al. [145] investigated the effectiveness of optical flow in recognizing full expressions, as well as micro-expressions, in near-frontal images. As far we know, only one work [146] has explored optical flow for FER with minimal annotations i.e., sequence-level labels.
@@
(g) Auxiliary Spaces for label Uncertainty Exploring the auxiliary label space of relevant tasks has recently gained much attention for dealing with label uncertainty. The idea is to leverage the information available in the auxiliary label space of related but distinct tasks to refine the label uncertainty of the main FER task. Chen et al. [147] explored the labels of AU recognition and facial landmark detection to effectively learn the label distributions based on the assumption that neighboring facial images in the auxiliary label space should have similar expression distributions. Le et al. [148] showed that leveraging the neighborhood information in the valence-arousal space helps to obtain more semantic label distributions, and further improves system performance. Unlike the above two approaches, Liu et al. [149] exploited the auxiliary tasks of AU detection and valence-arousal measurement to learn to relabel uncertain samples and mitigate the class imbalance respectively. Therefore, leveraging the auxiliary label space is a promising line of research to deal with the problem of label uncertainty, and can also be extended to deal with limited annotations.
@@
(h) Few Shot Learning Few-Shot Learning (FSL) is another emerging machine learning paradigm that deals with the problem of learning from limited annotated samples of novel classes. The basic idea of FSL is to generalize a pre-trained model to new classes by learning from a limited set of labeled samples of these classes (support set) [150]. Although FSL does not exactly correspond to any of the WSL scenarios, it overlaps somewhat with the framework of incomplete annotations, as both cases deal with the problem of limited accurate annotations. Recently, FSL has also gained attention in FER as a solution to the problem of novel classes for compound emotion recognition [151, 127, 152]. Another line of research with FSL for FER is to adapt to new subjects as people express their emotions differently from each other [153]. Most of the existing approaches pertinent to FSL for FER have been evaluated in cross-domain settings, where the target domain may have few labeled samples of novel classes [151, 152, 127]. Xinyi et al. [151] proposed a novel emotion-guided similarity network to deal with the limited annotations of compound expressions, where the pre-trained model is generalized to limited samples of novel classes (support set) of compound emotions. Zou et al. [152] also addressed the problem of compound FER in a cross-domain setting using a cascaded decomposition network for FSL. Chen et al. [127] explored self-supervised vision transformers by jointly pertaining with multiple pretext tasks, and leveraged FSL to train the deep model with fewer labeled samples. Besides these few papers on compound emotion recognition, FSL has not been explored for AU recognition. Therefore, the use of FSL techniques is a promising line of research to deal with limited labeled samples for FER.
VI Conclusion